# Absolute Calibration of Cluster Mira Variables to Provide a New Anchor for the Hubble Constant Determination

### v1

@@ Abstract
Mira variables in globular clusters can provide an accurate and precise absolute calibration of their period-luminosity relations (PLRs) to independently anchor the cosmic distance scale and determine the Hubble constant. We present homogeneous near-infrared () time-series photometric observations of a sample of 55 candidate long-period variables in 18 globular clusters covering a wide metallicity range ( dex). The Gaia proper motions, long-period variability information, and optical-infrared colors are used to identify 41 Oxygen-rich Miras as members of the globular clusters. Mean luminosities of Miras in the bands are independently calibrated using the recommended distances and mean parallaxes to their host clusters. Cluster Mira PLRs exhibit scatter comparable to the Large Magellanic Cloud (LMC) variables and do not show any dependence on iron-abundance for a wide range of metallicities. We establish the accuracy of cluster Miras as independent anchors by determining a distance modulus to the LMC, mag, in agreement with the 1.2% precise geometric distance. Our -band photometry is transformed to derive Hubble Space Telescope F160W PLR for cluster Miras providing a three-anchor baseline with the LMC and NGC 4258. We employ three-anchor solution to determine distances to two type Ia supernovae host galaxies, NGC 1559 ( mag) and M101 ( mag), and provide a measurement of the Hubble constant, km s-1 Mpc-1. Similar to Cepheids, our independent baseline solution results in a local determination that is systematically larger than its inference from the early universe probes, further supporting the ongoing Hubble tension.
@@ 1 Introduction
From the earliest studies on NIR observations of Miras (e.g., I. S. Glass & T. Lloyd Evans, 1981; J. W. Menzies & P. A. Whitelock, 1985; M. W. Feast et al., 1989), it was well-known that these variables follow a tight period-luminosity relation (PLR) with a typical scatter of mag using single or a few epochs of measurements. Later, the studies of LPVs in microlensing surveys led to the discovery of multiple PLRs for giants (P. R. Wood, 2000; M. R. L. Cioni et al., 2001; Y. Ita et al., 2004). With modern time-domain photometric surveys and infrared detectors, Mira PLRs, in particular in the Large Magellanic Cloud (LMC), have been obtained with accuracy and precision comparable to the classical Cepheids in the band ( mag, Y. Ita & N. Matsunaga, 2011; W. Yuan et al., 2017). This has led to Miras being calibrated at a high enough precision to determine Mira-based distances to SNe Ia host galaxies providing an independent determination (C. D. Huang et al., 2024).
Currently, Miras in the LMC and maser-host galaxy NGC 4258 play the crucial role of calibrating their PLR using geometric distances to these galaxies. The absolute calibration of Miras in our own Galaxy is difficult due to uncertainties associated with parallaxes of these bright, red stars with extended atmospheres and large amplitude variations, and lack of homogeneous reddening and metallicity estimates. J. L. Sanders (2023) used Gaia data release 3 (DR3) parallaxes of Milky Way field variables to calibrate the Mira distance scale. However, large parallax uncertainties coupled with random-epoch NIR observations from Two Micron All Sky Survey (2MASS, M. F. Skrutskie et al., 2006) results in large scatter in the Mira PLR, making those unreliable for a precise determination (also see discussion in C. D. Huang et al., 2024). Mira variables in star clusters can potentially provide a more accurate and precise calibration of their PLRs due to availability of independent distances and improvements in the mean parallax from thousands of cluster member stars (e.g. for Cepheids in A. G. Riess et al., 2022b). This work presents homogeneous time-series photometry of Mira variables in globular clusters and use their luminosity calibrations to determine a precise baseline value of together with LMC and NGC 4258 as independent anchors.
@@ 2 The Data
To obtain accurate periods, we extended the time-baseline of NIR observations of our cluster LPVs by compiling additional data from J. A. Frogel et al. (1981) for four NGC 104 stars and from J. W. Menzies & P. A. Whitelock (1985) for 16 variables. These literature datasets were homogenized by M. Feast et al. (2002) in the SAAO system, and were transformed to the 2MASS system using transformations from C. Koen et al. (2007).
The sample of 55 IRSF LPVs was cross-matched with the Gaia DR3 astrometry and the LPV catalogs (Gaia Collaboration et al., 2023; T. Lebzelter et al., 2023). While all IRSF LPVs have Gaia astrometry, only 44 (of 55) variables are listed in the Gaia LPV catalog, for which we compile preliminary periods, variability amplitude, and C-type star flag, if available. The sub-classification of all 55 IRSF LPVs in Mira (M), semi-regular (SR) variables was obtained from the C. M. Clement et al. (2001) catalog of variable stars, and uncertain classifications were flagged as M/SR subtype.
Fig. 1 displays spatial distribution, proper motions, and optical-NIR color-magnitude diagrams of LPVs in our sample in two representative clusters, N6441 and N104. N6441 has the largest number of LPVs in our sample (Ter 5 also has 8 LPVs) and is located in the bulge with relatively high reddening and crowding. N104 represents the cluster with the second largest number of five LPVs, but has least reddening in the GC sample. All sources that are within from the cluster center and have both Gaia and 2MASS photometry are shown. These properties of LPVs displayed in Fig. 1 for all clusters are used in Section 3 to identify cluster members and O-rich Mira variables.
@@
Note. — The cluster names are truncated throughout the text (see Table 2 for full names). The ‘Flag’ Yes/No refers to LPVs used in the initial PLR analysis. The remarks column provides reasons for exclusion of LPVs from the PLR fits. aThe band magnitude for Pal6 V1 is taken from 2MASS.
@@ 2.1 Period and mean-magnitude determinations
The IRSF light curves of 55 LPVs are shown in the appendix A (Figures 7–9). They are phased using their long-term periods. LPVs are not strictly periodic and our goal is to obtain their true mean magnitudes corresponding to the long-term period. Therefore, we fitted a sinusoidal template to the phased light curve in each filter assuming negligible phase lag between three NIR filters. The intensity-averaged mean magnitudes and peak-to-peak amplitudes were obtained from the best-fitted template light curves in filters. Table 1 lists the pulsation properties of all 55 IRSF LPVs in our sample. The photometric properties of 40 field Miras in the Milky Way are also provided in the appendix B (Table 7).
@@ 2.2 Distances, reddening, and metallicity
The LPV sample was cross-matched with the globular cluster distances catalog provided by H. Baumgardt & E. Vasiliev (2021). There are 40 LPVs in 15 clusters that are within 10 kpc distance, and the median uncertainty on distance for the entire sample is . The mean parallaxes of globular clusters were also obtained from E. Vasiliev & H. Baumgardt (2021) which were computed taking into account spatially correlated systematic errors after incorporating parallax-corrections by L. Lindegren et al. (2021). Since several clusters in our sample are in the bulge, where crowding and extinction uncertainties dominate, the average uncertainty on the parallaxes is . Furthermore, E. Vasiliev & H. Baumgardt (2021) also suggested that parallax uncertainties are underestimated by and the parallaxes are overestimated by mas for mag stars after applying L. Lindegren et al. (2021) correction. The parallax over-correction has also been noted for several other types of pulsating stars like RR Lyrae and Cepheids (A. Bhardwaj et al., 2021; A. G. Riess et al., 2021). We also correct parallaxes for field Miras for the parallax zero-point correction from L. Lindegren et al. (2021).
@@ 3.2 Mira Period-Luminosity relations
where, is the absolute magnitude in bands, and and represent the zero-point and slope, respectively, at a given wavelength. We obtained a preliminary linear regression fit to our sample of 55 LPVs after iteratively removing outliers. We found 8 LPVs (Lynga7 V1, Pal6 V1, N6441 V25, N6553 V5, N6637 V3, Ter5 V1, Ter5 V9, Ter12 V2) exhibiting residuals larger than 0.5 mag in all three filters. Seven of these eight LPVs have been identified as outliers in the previous subsection, confirming their exclusion as non-cluster members or potential C-rich stars or highly reddened O-rich Miras. The residuals of N6553 V5 exceed 0.8 mag in NIR PLRs, suggesting its period of 119.2, estimated from sparsely sampled light curves, may not be correct. This variable was not classified as LPV in the Gaia and does not have any period determination in the C. M. Clement et al. (2001) catalog. Therefore, N6553 V5 may not be a LPV and is excluded from the subsequent analysis.
@@
To investigate the impact of reddening corrections, we also derive PW relations in NIR filters, defined as: , , and , based on the E. L. Fitzpatrick (1999) extinction law (see Section 2.2). The slope and zero-points of the period-Wesenheit relations (PWRs) are consistent between the LMC and cluster Miras, except for , in which the slopes are statistically different. Note that PWRs exhibit comparably large scatter in both the LMC and globular cluster. These Wesenheit color coefficients are different to those adopted by W. Yuan et al. (2017), and are not necessarily appropriate as the interstellar extinction law may differ to the circumstellar dust extinction in Mira variables.
@@
Since our sample of cluster Miras cover a wide metallicity range ( [Fe/H] ), we investigated residual variations and found no significant trend as a function of metallicity. To quantify possible metallicity effects, we also solved for an additional metallicity term in the PLRs and PWRs for Mira variables in clusters and found a null or negligible dependence ( mag/dex) on iron abundance in bands. We also separated O-rich Mira sample to metal-rich ([Fe/H] dex, 21 stars) and metal-poor ([Fe/H] dex) samples and derived their separate PLRs. While the slopes are similar within for the two subsample, the zero-points are statistically consistent within in all-three bands. We conclude there is no significant metallicity effect on the PLRs for cluster Mira variables. Similar null or small effect of metallicity has also been noted in previous studies on Miras in globular clusters and LMC, and AGB stars in nearby metal-poor galaxies (M. Rejkuba, 2004; P. A. Whitelock et al., 2008; A. Bhardwaj et al., 2019; S. R. Goldman et al., 2019). However, the majority of LPVs in our sample belong to clusters with a narrow range of metallicity, [Fe/H], which can limit a proper quantification of metallicity dependence of PLRs for cluster Mira variables.
@@ 3.3 Astrometry-based luminosity fits
The absolute calibration of Mira PLRs listed in Table 3 are based on recommended distances to the clusters by H. Baumgardt & E. Vasiliev (2021). However, homogeneous Gaia DR3 parallax-based distances are not available for the entire sample of clusters in this catalog. Therefore, we also use mean parallaxes for globular clusters to calibrate NIR PLRs for Mira variables (E. Vasiliev & H. Baumgardt, 2021). As mentioned in the Section 2.2, the average uncertainty on the mean parallaxes of clusters in our sample is . Therefore, the calibration is obtained using astrometry-based luminosity (ABL, M. W. Feast & R. M. Catchpole, 1997; F. Arenou & X. Luri, 1999) defined as:
@@
where is the extinction corrected apparent mean magnitude in . The results of the ABL fits are shown in Table 4 after iteratively removing the outliers. The slopes and zero-points obtained using the ABL fits are within of those listed in Table 3 in all cases, showing consistency with the PLR calibrations based on globular cluster distances. Nevertheless, the zero-points based on this approach systematically differ by 0.05/0.06/0.04 mag in , hinting an over-correction in cluster parallaxes as suggested by E. Vasiliev & H. Baumgardt (2021).
@@ 3.4 Distance to the LMC
We note that the LMC Mira sample was not corrected for interstellar reddening. Therefore, we estimate color-excess, , from the reddening maps of D. M. Skowron et al. (2021), and find a median value of mag at the location of O-rich Miras. The reddening value was converted to in D. J. Schlegel et al. (1998) scale using transformations from D. M. Skowron et al. (2021). Mira magnitudes were corrected for extinction using the E. L. Fitzpatrick (1999) reddening law as discussed in Section 2.2. The results of fitting the fixed slopes in are shown in Table 5, where the LMC distance agrees well with the most precise value of from late-type eclipsing binaries (G. Pietrzyński et al., 2019). The uncertainty on the LMC distance modulus includes the systematics of 0.029 mag in calibrator cluster distances. The -band distance is smaller by than those derived from and band PLRs. The average true distance modulus of mag using bands is in very good agreement with the geometric distance to the LMC (G. Pietrzyński et al., 2019), thus establishing the use of Miras in globular clusters to constrain their accurate luminosity zero-points in the NIR bands.
@@ 4.1 Ground-based to HST F160W PLR
To complement the observations of Miras in SNe Ia host galaxies in HST F160W photometric system, C. D. Huang et al. (2018) derived a color transformation: . In the case of globular clusters, we used extinction corrected mean magnitudes in and bands transforming to F160W magnitudes. Using the same initial sample of 41 Miras as in Table 3, a PLR in the F160W filter was derived:
with a scatter of 0.15 mag for stars with periods smaller than 400 d. Fig. 4 displays this PLR in HST F160W filter. We do not find any trends in the residuals of the best-fitting relation as a function of periods of Miras, and with metallicity or reddening of their host clusters. The globular clusters cover a wide metallicity range , thus these results further show that cluster Mira PLRs are not sensitive to metallicity effects as previously suggested for Mira/AGB stars in several studies at infrared wavelengths (P. A. Whitelock et al., 2008; A. Bhardwaj et al., 2019; S. R. Goldman et al., 2019).

### v2

@@ Abstract
Mira variables in globular clusters can provide an accurate and precise absolute calibration of their period-luminosity relations (PLRs) to independently anchor the cosmic distance scale and determine the Hubble constant. We present homogeneous near-infrared () time-series photometric observations of a sample of 55 candidate long-period variables in 18 globular clusters covering a wide metallicity range ( dex). The Gaia proper motions, long-period variability information, and optical-infrared colors are used to identify 41 oxygen-rich Miras as members of the globular clusters. Mean luminosities of Miras in the bands are independently calibrated using the recommended distances and mean parallaxes to their host clusters. Cluster Mira PLRs exhibit scatter comparable to the Large Magellanic Cloud (LMC) variables and do not show any dependence on iron abundance for a wide range of metallicities. We establish the accuracy of cluster Miras as independent anchors by determining a distance modulus to the LMC, mag, in agreement with the 1.2% precise geometric distance. Our -band photometry is transformed to derive Hubble Space Telescope F160W PLR for cluster Miras providing a three-anchor baseline with the LMC and NGC 4258. We employ three-anchor solution to determine distances to two type Ia supernovae host galaxies, NGC 1559 ( mag) and M101 ( mag), and provide a measurement of the Hubble constant, km s-1 Mpc-1. Similar to Cepheids, our independent baseline solution results in a local determination that is systematically larger than its inference from the early universe probes, further supporting the ongoing Hubble tension.
@@ 1 Introduction
From the earliest studies on NIR observations of Miras (e.g., I. S. Glass & T. Lloyd Evans, 1981; J. W. Menzies & P. A. Whitelock, 1985; M. W. Feast et al., 1989), it was well known that these variables follow a tight period-luminosity relation (PLR) with a typical scatter of mag using single or a few epochs of measurements. Later, the studies of LPVs in microlensing surveys led to the discovery of multiple PLRs for giants (P. R. Wood, 2000; M. R. L. Cioni et al., 2001; Y. Ita et al., 2004). With modern time-domain photometric surveys and infrared detectors, Mira PLRs, in particular in the Large Magellanic Cloud (LMC), have been obtained with accuracy and precision comparable to the classical Cepheids in the band ( mag, Y. Ita & N. Matsunaga, 2011; W. Yuan et al., 2017). This has led to Miras being calibrated at a high enough precision to determine Mira-based distances to SNe Ia host galaxies providing an independent determination (C. D. Huang et al., 2024).
Currently, Miras in the LMC and maser-host galaxy NGC 4258 play the crucial role of calibrating their PLR using geometric distances to these galaxies. The absolute calibration of Miras in our own Galaxy is difficult due to uncertainties associated with parallaxes of these bright, red stars with extended atmospheres and large amplitude variations, and lack of homogeneous reddening and metallicity estimates. J. L. Sanders (2023) used Gaia data release 3 (DR3) parallaxes of Milky Way field variables to calibrate the Mira distance scale. However, large parallax uncertainties coupled with random-epoch NIR observations from Two Micron All Sky Survey (2MASS, M. F. Skrutskie et al., 2006) result in large scatter in the Mira PLR, making those unreliable for a precise determination (also see discussion in C. D. Huang et al., 2024). Mira variables in star clusters can potentially provide a more accurate and precise calibration of their PLRs due to availability of independent distances and improvements in the mean parallax from thousands of cluster member stars (e.g. for Cepheids in A. G. Riess et al., 2022b). This work presents homogeneous time-series photometry of Mira variables in globular clusters and uses their luminosity calibrations to determine a precise baseline value of together with LMC and NGC 4258 as independent anchors.
@@ 2 The Data
To obtain accurate periods, we extended the time baseline of NIR observations of our cluster LPVs by compiling additional data from J. A. Frogel et al. (1981) for four NGC 104 stars and from J. W. Menzies & P. A. Whitelock (1985) for 16 variables. These literature datasets were homogenized by M. Feast et al. (2002) in the SAAO system, and were transformed to the 2MASS system using transformations from C. Koen et al. (2007).
The sample of 55 IRSF LPVs was crossmatched with the Gaia DR3 astrometry and the LPV catalogs (Gaia Collaboration et al., 2023; T. Lebzelter et al., 2023). While all IRSF LPVs have Gaia astrometry, only 44 (of 55) variables are listed in the Gaia LPV catalog, for which we compile preliminary periods, variability amplitude, and C-type star flag, if available. The subclassification of all 55 IRSF LPVs in Mira (M), semi-regular (SR) variables was obtained from the C. M. Clement et al. (2001) catalog of variable stars, and uncertain classifications were flagged as M/SR subtype.
Fig. 1 displays spatial distribution, proper motions, and optical-NIR color-magnitude diagrams of LPVs in our sample in two representative clusters, N6441 and N104. N6441 has the largest number of LPVs in our sample (Ter 5 also has 8 LPVs) and is located in the bulge with relatively high reddening and crowding. N104 represents the cluster with the second largest number of five LPVs, but has least reddening in the Globular Cluster (GC) sample. All sources that are within from the cluster center and have both Gaia and 2MASS photometry are shown. These properties of LPVs displayed in Fig. 1 for all clusters are used in Section 3 to identify cluster members and O-rich Mira variables.
@@
Note. — The cluster names are truncated throughout the text (see Table 2 for full names). The ‘Flag’ Yes/No refers to LPVs used in the initial PLR analysis. The remarks column provides reasons for exclusion of LPVs from the PLR fits. aThe band magnitude for Pal6 V1 is taken from 2MASS (M. F. Skrutskie et al., 2006).
@@ 2.1 Period and mean-magnitude determinations
The IRSF light curves of 55 LPVs are shown in Appendix A (Figures 7–9). They are phased using their long-term periods. LPVs are not strictly periodic and our goal is to obtain their true mean magnitudes corresponding to the long-term period. Therefore, we fitted a sinusoidal template to the phased light curve in each filter assuming negligible phase lag between three NIR filters. The intensity-averaged mean magnitudes and peak-to-peak amplitudes were obtained from the best-fitted template light curves in filters. Table 1 lists the pulsation properties of all 55 IRSF LPVs in our sample. The photometric properties of 40 field Miras in the Milky Way are also provided in Appendix B (Table 7).
@@ 2.2 Distances, reddening, and metallicity
The LPV sample was crossmatched with the globular cluster distances catalog provided by H. Baumgardt & E. Vasiliev (2021). There are 40 LPVs in 15 clusters that are within 10 kpc distance, and the median uncertainty on distance for the entire sample is . The mean parallaxes of globular clusters were also obtained from E. Vasiliev & H. Baumgardt (2021) which were computed taking into account spatially correlated systematic errors after incorporating parallax corrections by L. Lindegren et al. (2021). Since several clusters in our sample are in the bulge, where crowding and extinction uncertainties dominate, the average uncertainty on the parallaxes is . Furthermore, E. Vasiliev & H. Baumgardt (2021) also suggested that parallax uncertainties are underestimated by and the parallaxes are overestimated by mas for mag stars after applying L. Lindegren et al. (2021) correction. The parallax overcorrection has also been noted for several other types of pulsating stars like RR Lyrae and Cepheids (A. Bhardwaj et al., 2021; A. G. Riess et al., 2021). We also correct parallaxes for field Miras for the parallax zero-point correction from L. Lindegren et al. (2021).
@@ 3.2 Mira Period-Luminosity relations
where, is the absolute magnitude in bands, and and represent the zero-point and slope, respectively, at a given wavelength. We obtained a preliminary linear regression fit to our sample of 55 LPVs after iteratively removing outliers. We found 8 LPVs (Lynga7 V1, Pal6 V1, N6441 V25, N6553 V5, N6637 V3, Ter5 V1, Ter5 V9, Ter12 V2) exhibiting residuals larger than 0.5 mag in all three filters. Seven of these eight LPVs have been identified as outliers in the previous subsection, confirming their exclusion as non-cluster members or potential C-rich stars or highly reddened O-rich Miras. The residuals of N6553 V5 exceed 0.8 mag in NIR PLRs, suggesting its period of 119.2, estimated from sparsely sampled light curves, may not be correct. This variable was not classified as LPV in the Gaia and does not have any period determination in the C. M. Clement et al. (2001) catalog. Therefore, N6553 V5 may not be an LPV and is excluded from the subsequent analysis.
@@
To investigate the impact of reddening corrections, we also derive PW relations in NIR filters, defined as: , , and , based on the E. L. Fitzpatrick (1999) extinction law (see Section 2.2). The slope and zero-points of the period-Wesenheit relations (PWRs) are consistent between the LMC and cluster Miras, except for , in which the slopes are statistically different. Note that PWRs exhibit comparably large scatter in both the LMC and globular cluster. These Wesenheit color coefficients are different from those adopted by W. Yuan et al. (2017), and are not necessarily appropriate as the interstellar extinction law may differ from the circumstellar dust extinction in Mira variables.
@@
Since our sample of cluster Miras covers a wide metallicity range ( [Fe/H] ), we investigated residual variations and found no significant trend as a function of metallicity. To quantify possible metallicity effects, we also solved for an additional metallicity term in the PLRs and PWRs for Mira variables in clusters and found a null or negligible dependence ( mag/dex) on iron abundance in bands. We also separated the O-rich Mira sample into the metal-rich ([Fe/H] dex, 21 stars) and metal-poor ([Fe/H] dex) samples and derived their separate PLRs. While the slopes are similar within for the two subsamples, the zero-points are statistically consistent within in all three bands. We conclude there is no significant metallicity effect on the PLRs for cluster Mira variables. A similar null or small effect of metallicity has also been noted in previous studies on Miras in globular clusters and LMC, and AGB stars in nearby metal-poor galaxies (M. Rejkuba, 2004; P. A. Whitelock et al., 2008; A. Bhardwaj et al., 2019; S. R. Goldman et al., 2019). However, the majority of LPVs in our sample belong to clusters with a narrow range of metallicity, [Fe/H], which can limit a proper quantification of metallicity dependence of PLRs for cluster Mira variables.
@@ 3.3 Astrometry-based luminosity fits
The absolute calibration of Mira PLRs listed in Table 3 is based on recommended distances to the clusters by H. Baumgardt & E. Vasiliev (2021). However, homogeneous Gaia DR3 parallax-based distances are not available for the entire sample of clusters in this catalog. Therefore, we also use mean parallaxes for globular clusters to calibrate NIR PLRs for Mira variables (E. Vasiliev & H. Baumgardt, 2021). As mentioned in Section 2.2, the average uncertainty on the mean parallaxes of clusters in our sample is . Therefore, the calibration is obtained using astrometry-based luminosity (ABL, M. W. Feast & R. M. Catchpole, 1997; F. Arenou & X. Luri, 1999) defined as:
@@
where is the extinction corrected apparent mean magnitude in . The results of the ABL fits are shown in Table 4 after iteratively removing the outliers. The slopes and zero-points obtained using the ABL fits are within of those listed in Table 3 in all cases, showing consistency with the PLR calibrations based on globular cluster distances. Nevertheless, the zero-points based on this approach systematically differ by 0.05/0.06/0.04 mag in , hinting an overcorrection in cluster parallaxes as suggested by E. Vasiliev & H. Baumgardt (2021).
@@ 3.4 Distance to the LMC
We note that the LMC Mira sample was not corrected for interstellar reddening. Therefore, we estimate color-excess, , from the reddening maps of D. M. Skowron et al. (2021), and find a median value of mag at the location of O-rich Miras. The reddening value was converted to in D. J. Schlegel et al. (1998) scale using transformations from D. M. Skowron et al. (2021). Mira magnitudes were corrected for extinction using the E. L. Fitzpatrick (1999) reddening law as discussed in Section 2.2. The results of fitting the fixed slopes in are shown in Table 5, where the LMC distance agrees well with the most precise value of from late-type eclipsing binaries (G. Pietrzyński et al., 2019). The uncertainty on the LMC distance modulus includes the systematics of 0.029 mag in calibrator cluster distances. The -band distance is smaller by than that derived from and band PLRs. The average true distance modulus of mag using bands is in very good agreement with the geometric distance to the LMC (G. Pietrzyński et al., 2019), thus establishing the use of Miras in globular clusters to constrain their accurate luminosity zero-points in the NIR bands.
@@ 4.1 Ground-based to HST F160W PLR
To complement the observations of Miras in SNe Ia host galaxies in the HST F160W photometric system, C. D. Huang et al. (2018) derived a color transformation: . In the case of globular clusters, we used extinction corrected mean magnitudes in and bands transforming to F160W magnitudes. Using the same initial sample of 41 Miras as in Table 3, a PLR in the F160W filter was derived:
with a scatter of 0.15 mag for stars with periods smaller than 400 d. Fig. 4 displays this PLR in the HST F160W filter. We do not find any trends in the residuals of the best-fitting relation as a function of periods of Miras, and with metallicity or reddening of their host clusters. The globular clusters cover a wide metallicity range , thus these results further show that cluster Mira PLRs are not sensitive to metallicity effects as previously suggested for Mira/AGB stars in several studies at infrared wavelengths (P. A. Whitelock et al., 2008; A. Bhardwaj et al., 2019; S. R. Goldman et al., 2019).